The product of two 2-digit numbers is 1665. The product of their units digit is
35 and that of their tens digit is 12. Find the two numbers.

Answer by Edwin McCravy(20060)

Instead of doing it for you, I'll do another
one exactly like it and explain every step, 
so you can do yours exactly the same way. 
I will do and explain this one instead:

The product of two 2-digit numbers is 2632.
The product of their units digit is 32 and
that of their tens digit is 18. Find the two
numbers.
Start with this:

...the product of their units digit is 32...
The only two digits that have product 32 are 4 and 8
So their units digits are 4 and 8

...and that of tens digit is 18...
So their tens digits could be
3 and 6 or 2 and 9

The possibilities are:

34 and 68
64 and 38
24 and 98
94 and 28

The product of two 2-digit numbers is 2632...
So we multiply them to find out:

34×68 = 2312
64×38 = 2432
24×98 = 2352
94×28 = 2632

So they are 94 and 28.

Now do yours exactly the same way.

Edwin